OpenStudy (anonymous):

Easy calculus question: if you had something like say a double integral integral[integral (x^2+y^2)/(xy) dy]dx why can't you set x^2 + y^2 = 1?

OpenStudy (anonymous):

using the cos^2x + sin^2y identity

OpenStudy (anonymous):

\[\int\limits_{1}^{4}\int\limits_{1}^{2}(x^2+y^2)/xy)dydx\]

OpenStudy (anonymous):

why can't the x^2 + y^2 be set = 1

OpenStudy (amistre64):

i think you are confusing a few things

OpenStudy (amistre64):

\[\int\int \frac{x^2+y^2}{xy} dy.dx\] \[\int\int \frac{x^2}{xy}+\frac{y^2}{xy} dy.dx\] \[\int\int \frac{x}{y}+\frac{y}{x} dy.dx\] \[\int\int xln(y)+\frac{y^2}{2x} \ dx\]etc... if that equates to your idea, then try it out

OpenStudy (anonymous):

ah sorry didn't mean to make you write all of that out that's actually what the original equation was, i just moved evertying around to try something out

OpenStudy (anonymous):

was just kinda wondering when you can or can't use the x^2+y^2 = cos[x]^2 + sin[y]^2 = 1 identities

OpenStudy (amistre64):

oh, i wrote it for the practice:)

OpenStudy (anonymous):

just figured that since we are using double integrals, they are equations with x and y, so i was thinking you could use identities to maybe simplify things

OpenStudy (anonymous):

:) heh

OpenStudy (amistre64):

(..)^2 is not generally equativalent to the function trig^2(...)

OpenStudy (anonymous):

You can use that identity to simplify integrals but you have to be carefull.

OpenStudy (anonymous):

yeah, the integral didn't come out right in mathematica. Or, could you use it but you have to change the limits of integration?

OpenStudy (anonymous):

\[x^2+y^2\neq \cos^2x+\sin^2x\]

OpenStudy (anonymous):

ah right but x^2/r^2 + y^2/r^2 = cos[x]^2 + sin[y]^2 = r^2 right?

OpenStudy (turingtest):

no

OpenStudy (turingtest):

\[x=r\cos\theta\]\[y=r\sin\theta\]\[x^2+y^2=r^2\cos^2\theta+r^2\sin^2\theta=r^2\]I'm not sure what you are trying to get out of that

OpenStudy (anonymous):

er yeah, sorry messed that up. it would still work out though right?? because x^2/1 + y^2/1 = (1)^2 cos^2 + sin^2 = 1^2 blehhh i think i'm mixing this all up

OpenStudy (anonymous):

Don't worry.

OpenStudy (turingtest):

you are thinking something like\[x^2+y^2=r^2\implies{x^2+y^2\over r^2}=1\]perhaps? That would totally screw up your coordinate system though, to put that in the integral

OpenStudy (turingtest):

thta's some crazy cartesian-polar hybrid

OpenStudy (anonymous):

hahaha

OpenStudy (amistre64):

but its good to see you thinking outside the mathical box

OpenStudy (anonymous):

ah i haven't read the chapter on polar stuff yet, so maybe i'll remedy things in the coming week :)

OpenStudy (anonymous):

thank you all for the help!

OpenStudy (turingtest):

welcome :)